OpenStudy (sloppycanada):

Write 15 degrees 26' 29" using decimal degrees. Round your answer to the nearest thousandth. I don't want someone to do this for me, please explain it to me.

OpenStudy (prizzyjade):

that is something you need to use a calculator

OpenStudy (prizzyjade):

scientific calculator to be exact

OpenStudy (sloppycanada):

@jim_thompson5910 When you get the chance...

jimthompson5910 (jim_thompson5910):

the `26'` means `26 minutes` or `26 arcminutes` the `29" ` means `29 seconds` or `29 arcseconds`

jimthompson5910 (jim_thompson5910):

Rules: 1 degree = 60 arcminutes 1 arcminute = 60 arcseconds

jimthompson5910 (jim_thompson5910):

sound familiar?

OpenStudy (sloppycanada):

Yes.

jimthompson5910 (jim_thompson5910):

using those rules, we can write `26 arcminutes` as `(26/60) degrees` 60 arcminutes will give you a full degree. So 26 arcminutes is going to be a fraction (smaller than 1) of a degree

jimthompson5910 (jim_thompson5910):

1 degree = 60 arcminutes 1 arcminute = 60 arcseconds means 1 degree = 60*60 = 3600 arc seconds so `29 arcseconds = (29/3600) degrees`

OpenStudy (sloppycanada):

Do I divide?

OpenStudy (sloppycanada):

That would be .008

jimthompson5910 (jim_thompson5910):

so, 15 degrees 26' 29" = (15 degrees) + (26/60 degrees) + (29/3600 degrees) 15 degrees 26' 29" = (15 + 26/60 + 29/3600) degrees 15 degrees 26' 29" = ??? degrees

OpenStudy (sloppycanada):

15.441

jimthompson5910 (jim_thompson5910):

yeah I'm getting 15.4413888888889

OpenStudy (sloppycanada):

If I were to round that to the nearest thousandth, it's be 15.440

jimthompson5910 (jim_thompson5910):

15.441 actually

OpenStudy (sloppycanada):

Alright, the next thing asks me to change the 29 to 9. so it'd be 15 degrees 26' and 9"

OpenStudy (sloppycanada):

Would that just be 15.436?

jimthompson5910 (jim_thompson5910):

looks good 15 + 26/60 + 9/3600 = 15.4358333333333 ---> 15.436

OpenStudy (sloppycanada):

45 degrees 9' 30"

OpenStudy (sloppycanada):

45 + 9/60 + 30/3600 = 45.158

jimthompson5910 (jim_thompson5910):

looks good \[\color{red}{45} \text{ degrees}, \color{blue}{9} \text{ minutes}, \color{green}{30} \text{ seconds} = (\color{red}{45} + \color{blue}{9}/60 + \color{green}{30}/3600) \text{degrees}\]